by intersection "x" and "y" graphic which are the intersection points of
2x -y =3 : x + y=6
b) x +3y = 18 ; 2x+ y = 11

Answer by stanbon(75887)

which are the intersection points of
2x -y =3
If x = 0, y = -3
If y = 0, x = 3/2
Plot (0,-3) and (3/2,0) and draw a line thru them.
-----------
x + y=6
If x = 0, y = 6
If y = 0, x = 6
Plot (0,6) and (6,0) and draw a line thru them.
----
Find the intersection::
y = 2x-3
y = 6-x
----
2x-3 = 6-x
3x = 9
x = 3, y = 3
------
b) x +3y = 18 ; 2x+ y = 11
y = (-1/3)x + 6
y = -2x+11
----
Solve for "x"::
(-1/3)x + 6 = -2x+11
(5/3)x = 5
x = 3 ; then y = -2x+11 = 5
-----------
Cheers,
Stan H.
-----------